# Sketch-guided Cage-based 3D Gaussian Splatting Deformation

### v2

@@ 3.5 Implementation details
The cage is generated automatically by first extracting a mesh from the GS model using the coarse stage of the SuGaR [8] method. This mesh is very large, sometimes it is not a closed manifold and it has many fold-overs so it would not be suited for a cage. Therefore, we compute a triangulated offset surface using a function from Libigl [14] that applies marching cubes on a grid of signed distance values from the input triangle mesh. The resulting mesh is a close manifold triangular mesh with an adjustable number of vertices depending on the model size and level of detail desired. We used StableDiffusion-XL as the diffusion model of the ControlNet. For each deformation, we optimized for $2000$ iterations, with a learning rate of $0.002$, and optimized by Adam optimizer [18]. Except for the reference view, $4$ random views were sampled for the 3D-aware SDS. The diffusion model used in 3D-aware SDS is zero-1-to-3 XL [22]. All experiments were run on a single Nvidia RTX A6000 GPU. The running time was related to the number of Gaussian splats in different scenes. For the scene with less than 100k Gaussian splats, the running time is less than 10 minutes. More efficient diffusion models is a growing area of research and our method would be directly accelerated by the many methods that are being developed [32, 11].
@@ 4.1 Results
Figure 7: Animate static 3D GS by applying our method to key frames. Users can draw a few sketches for the key frames and apply our method to each key frame. The animation was generated by interpolating the deformed cages between key frames.
@@
We tested our method on various 3D objects, from human-made objects to animals and humans, as shown in Figure 1, 5, and 9. It shows that our method can deform the objects precisely with the guidance of the sketch and produce natural-looking results. These 3D shapes were collected from the SketchFab and TurboSquid and transferred into a 3D GS scene. The 3D shapes were originally mesh and converted into 3D GS by training from 100 sampling images from random views of the shape.
Moreover, we also explored the ability of our method to deform large-scale real-world captured data. We used the UAV-captured dataset of a great Oratory and a Clock Tower to reconstruct the 3D GS scene respectively. As shown in Figure 6.
As one of the important applications of 3D deformation is animation, we also conducted experiments of animating static 3D GS by our method. As shown in Figure 1 (C), 7, and the accompanying video, the user can provide multiple input sketches as some key-frame sketches in the animation sequence. Then, by running our method for each key frame and interpolating between the deformed cage of key frames, we can get animations of the static 3D GS.
Table 1: Quantitative comparison: we report the relative CLIP-IQA score [33], which is calculated as the CLIP-IQA score of the undeformed renderings divided by the CLIP-IQA score of the deformed renderings.
@@
Relative CLIP-IQA 75.32% 99.03%
@@
We compared our Cage-based deformation for 3D GS with the SOTA method MLP+HexPlane [28]. For comparison, we deform the source 3D GS using our pipeline and render the sketch view of the deformed 3D GS as the reference image of the MLP+Hexplane. Thus, except for the difference in the deformation method, MLP+Hexplane also has extra RGB image guidance in the sketch view when our pipeline is only guided by silhouette. Even in that case, our pipeline can deform the 3D GS with much higher fidelity than using MLP+Hexplane. The qualitative comparison was shown in Figure 5, because of the good space continuity of the cage deformation, the local detail features can be fully preserved, e.g. the pattern on the statue, and the teeth of the dinosaur. Though the Hexplane was applied as a geometric regularizer, the detailed features can still be destroyed and cause fuzzy renderings when using MLP+Hexplane. We also quantitatively compare the visual quality of the deformed shapes based on image quality. We use the CLIP-IQA [33], a metric measuring the image quality based on the CLIP score, to evaluate 8 deformed results. For every deformed 3D GS, 8 views were rendered as the images to be evaluated. Since our task is to measure how the image quality was changed after the deformation process, we report the relative CLIP-IQA score, which is calculated as the CLIP-IQA score of the undeformed renderings divided by the CLIP-IQA score of the deformed renderings. As shown in table 1, the average decrease of the CLIP-IQA score for the MLP+Hexplane method is $24.68\%$, however, our method almost keeps the same CLIP-IQA score as original renderings with less than $1\%$ decreasing.
@@ 4.2 Ablation
We tested the effects of two important components in our method, decomposed NJF and 3D-aware SDS. We explored the difference of optimizing directly on the cage vertices, via NJF and our proposed decomposed NJF. As shown in Figure 8, optimizing directly on the cage vertices can lead to undesired fuzzy rendering which is caused by entanglement of the cage during the optimization. As shown in Figure 8 (B), the detailed feather of the statue and the door on the side of the clock tower were destroyed. Applying NJF can solve this problem since it works as a geometry regularizer for the cage. However, it can be too strong to fit the sketched silhouette precisely. As shown in Figure 8 (C), the wing of the statue is not fitted on the end and the clock tower is not bent enough. When using decomposed NJF as stated in section 5, we can obtain a more efficient optimization which leads to a lower $\mathcal{L}_{sil}$ (average of $49.36\%$ decreasing in these two examples), but with same rendering visual quality compared to NJF (Figure 8 A).
We also explored the effect of 3D-aware SDS in our pipeline. As shown in Figure 9, the 3D-aware SDS plays a critical role in preserving undesirable deformation in the views except for the sketch view in the pipeline. For instance, when the pistol was deformed only by $\mathcal{L}_{sil}$, the barrel was bent, which can be fixed by adding the 3D-aware SDS.

### v3

@@ 3.5 Implementation details
The cage is generated automatically by first extracting a mesh from the GS model using the coarse stage of the SuGaR [guedon2024sugar] method. This mesh is very large, sometimes it is not a closed manifold and it has many fold-overs so it would not be suited for a cage. Therefore, we compute a triangulated offset surface using a function from Libigl [jacobson2013libigl] that applies marching cubes on a grid of signed distance values from the input triangle mesh. The resulting mesh is a close manifold triangular mesh with an adjustable number of vertices depending on the model size and level of detail desired. We used StableDiffusion-XL as the diffusion model of the ControlNet. For each deformation, we optimized for $2000$ iterations, with a learning rate of $0.002$, and optimized by Adam optimizer [kingma2014adam]. Except for the reference view, $4$ random views were sampled for the 3D-aware SDS. The diffusion model used in 3D-aware SDS is zero-1-to-3 XL [liu2023zero]. All experiments were run on a single Nvidia RTX A6000 GPU. The running time was related to the number of Gaussian splats and cage resolution in different scenes. For a scene with 268k gaussians and 376 cage vertices, the running time is 12 minutes. For a scene with 90k gaussians and 1343 cage vertices, the running time is 9 minutes. Thus, the running time is primarily influenced by the number of Gaussians. For large-scale scenes containing more Gaussians (e.g., 500k), the computation reaches a bottleneck, resulting in slower performance (around 40 minutes). More efficient diffusion models is a growing area of research and our method would be directly accelerated by the many methods that are being developed [starodubcev2024your, hong2024sfddm].
@@ 4.1 Results
Figure 6: The comparison of our method with mesh-binding method GaussianMesh [gao2024mesh]. From left to right: 1) the deformed proxy mesh obtained by our sketch-guided pipeline for GaussianMesh. 2) the deformed 3D GS by GaussianMesh. 3) deformation result obtained by our cage-based method. 4) undeformed original 3D GS.
@@
We tested our method on various 3D objects, from human-made objects to animals and humans, as shown in Figure 1, 5, and 9. It shows that our method can deform the objects precisely with the guidance of the sketch and produce natural-looking results. These 3D shapes were collected from the SketchFab and TurboSquid and transferred into a 3D GS scene. The 3D shapes were originally mesh and converted into 3D GS by training from 100 sampling images from random views of the shape. Moreover, we also explored the ability of our method to deform large-scale real-world captured data. We used the UAV-captured dataset of a great Oratory and a Clock Tower to reconstruct the 3D GS scene respectively. As shown in Figure 7.
As one of the important applications of 3D deformation is animation, we also conducted experiments of animating static 3D GS by our method. As shown in Figure 1(C), and the accompanying video, the user can provide multiple input sketches as some key-frame sketches in the animation sequence. Then, by running our method for each key frame and interpolating between the deformed cage of key frames, we can get animations of the static 3D GS.
Table 1: Quantitative comparison: we report the relative CLIP-IQA score [wang2023exploring], and relative Q-align score[wu2023q].
@@
Relative CLIP-IQA $\uparrow$ 75.32% 99.03%
Relative Q-align $\uparrow$ 78.11% 99.88%
@@
We first compared our Cage-based deformation for 3D GS with the SOTA method MLP+HexPlane [ren2023dreamgaussian4d]. For comparison, we deform the source 3D GS using our pipeline and render the sketch view of the deformed 3D GS as the reference image of the MLP+Hexplane. Thus, except for the difference in the deformation method, MLP+Hexplane also has extra RGB image guidance in the sketch view when our pipeline is only guided by silhouette. Even in that case, our pipeline can deform the 3D GS with much higher fidelity than using MLP+Hexplane. The qualitative comparison was shown in Figure 5, because of the good space continuity of the cage deformation, the local detail features can be fully preserved, e.g. the pattern on the statue, and the teeth of the dinosaur. Though the Hexplane was applied as a geometric regularizer, the detailed features can still be destroyed and cause fuzzy renderings when using MLP+Hexplane. We quantitatively compare the visual quality of the deformed shapes based on image quality. We use the CLIP-IQA [wang2023exploring], a metric measuring the image quality based on the CLIP score, and Q-align[wu2023q], an image quality assessment based on large multi-modality models (mPLUG-Owl2 [ye2024mplug]), to evaluate 8 deformed results. For every deformed 3D GS, 8 views were rendered as the images to be assessed. Since our task is to measure how the image quality was changed after the deformation process, we report the relative score of both CLIP-IQA and Q-align, which is calculated as the metric score of the undeformed renderings divided by the metric score of the deformed renderings. As shown in table 1, the average decreases of the CLIP-IQA score and Q-align score for the MLP+Hexplane method are $24.68\%$ and $21.89$ respectively, however, our method almost keeps the same CLIP-IQA and Q-align score as original renderings with less than $1\%$ decreasing. We also conducted a user study with 15 participants comparing the deformation fidelity to using MLP+HexPlane across 6 examples with instruction: which deformed result has better visual fidelity?. Among these 90 comparisons, our method gains $90.0\%$ preference from the participants.
We also include a comparison with mesh-binding GS (GaussianMesh [gao2024mesh]); however, a direct substitution of our cage-based method with their method is not straightforward, as their deformation approach is not implemented as a differentiable process. We perform the comparison with the following setup: 1) Run the mesh reconstruction method [instant-nsr-pl] to obtain a proxy mesh $M_{p}$, which is the base mesh to reconstruct the 3D GS. 2) Run GaussianMesh [gao2024mesh] reconstruction with input $M_{p}$ to get the 3D GS. 3) Run our pipeline as shown in figure 2 in which the 3D GS was replaced by $M_{p}$ and regularized by NJF to get a deformed mesh $M^{\prime}_{p}$. 4) Run GaussianMesh Deformation based on the deformed mesh $M^{\prime}_{p}$ to get the deformed 3D GS. The result is shown in figure 6. The deformed mesh exhibits severe artifacts, including entanglement and faces inversion, even with the NJF regularization. These artifacts are reflected in the final rendering of the 3D GS. Another drawback of the mesh-binding method is that the reconstruction is based on a reconstructed mesh obtained by other methods [instant-nsr-pl], which can fail in complex scenes. We show an example of a reconstructed mesh in the Supplementary material.
Additionally, we evaluated the silhouette adherence by the Intersection of Union (IoU) between the alpha rendering of 3D GS and the target silhouette from the sketch view. Among the examples, the average IoU before deformation is $0.654$ and after deformation is $0.901$.
@@ 4.2 Ablation
We tested the effects of two important components in our method, decomposed NJF and 3D-aware SDS. We explored the difference of optimizing directly on the cage vertices, via NJF and our proposed decomposed NJF. As shown in Figure 8, optimizing directly on the cage vertices can lead to undesired fuzzy rendering, which is caused by entanglement of the cage during the optimization. As shown in Figure 8 (B), the detailed feather of the statue and the door on the side of the clock tower were destroyed. Applying NJF can solve this problem since it works as a geometry regularizer for the cage. However, it can be too strong to fit the sketched silhouette precisely. As shown in Figure 8 (C), the wing of the statue is not fitted on the end and the clock tower is not bent enough. When using decomposed NJF as stated in section 5, we can obtain a more efficient optimization which leads to a lower $\mathcal{L}_{sil}$ (average of $49.36\%$ decreasing in these two examples), but with same rendering visual quality compared to NJF (Figure 8 A).
We also explored the effect of 3D-aware SDS in our pipeline. As shown in Figure 9, this plays a critical role in preserving undesirable deformation in the views except for the sketch view in the pipeline. For instance, when the pistol was deformed only by $\mathcal{L}_{sil}$, the barrel was bent, which can be fixed by adding the 3D-aware SDS.
4.3 Failure cases and mitigations
Some failures can occur from single view mismatches (e.g. mismatching of left and right legs in a side view). This can be addressed by slightly changing the view (See supplementary material)
@@ 6 Limitation and Future work
7 Acknowledgment
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), under funding reference numbers RGPIN-2021-03477 and RGPIN-2024-04605, as well as the support of Fonds de recherche du Québec – Nature et technologies (FRQNT), under funding reference number 365040.